Vancomycin is an antibiotic with volume of distribution of 0.65L/kg (this is an unusual
presentation of the volume of distribution).

If the target plasma concentration is 20mg/L was this the appropriate loading dose for the 70kg patient?

Why would you want to use a loading dose?
Answer #1

Volume of distribution of vancomycin = 0.65 L/kg

So total volume of distribution of the drug = 0.65* 70 = 45.5 L

Target plasma concentration = 20 mg/L

Loading dose = Target concentration * Volume of distribution

= 20 * 45.5 = 910 mg

Since vancomycin has a large volume of distribution, it indicates that the drug is sequestered in tissues.Hence the amount available in the plasma to act on the target site is less.Hence loading dose is required for rapid and more effective drug action.